OpenStudy (anonymous):

Find f(x) if y = f(x) satisfies \frac(dy)(dx) = 45 yx^(14) and the y -intercept of the curve y = f(x) is 3 .

OpenStudy (anonymous):

i don't know squat about differential equations, but my first guess would be to divide both sides by \(y\) and integrate

OpenStudy (anonymous):

\[\frac{y'}{y}=x^{14}\] \[\ln(y)=\frac{x^{15}}{15}\] etc

OpenStudy (anonymous):

oh i forgot the 45 sorry, \[\ln(y)=3x^{15}\]

OpenStudy (anonymous):

and so \(y=ce^{3x^{15}}\)

OpenStudy (anonymous):

but i would get a second opinion, because this is a stab looks good though, because you can check by differentiation

OpenStudy (psymon):

Yeah, nothing wrong with what ya did @satellite73

OpenStudy (anonymous):

I agree, but would c in this case be equal to 3? because is the yintercept.

OpenStudy (psymon):

We don't need c to equal anything in this case. We don't have any initial condition to straight out find C.

OpenStudy (anonymous):

ok, thanks!

OpenStudy (psymon):

Oh, there was an initial condition, my bad, haha.

OpenStudy (psymon):

Fail.

OpenStudy (psymon):

Nah, gotta get the full solution form first.

OpenStudy (loser66):

ok, you finish.

OpenStudy (psymon):

Might as well do the problem over, haha. \[\frac{ dy }{ dx }=45yx ^{14}\] \[\int\limits_{}^{}\frac{ dy }{ y }=45\int\limits_{}^{}x ^{14}\] \[lny = (45)\frac{ x ^{15} }{ 15 }+C \] \[e ^{lny}=e ^{3x ^{15}+C}\] \[y=e ^{C}*e ^{3x ^{15}}\] \[y=ce ^{3x ^{15}} \] y-intercepts occur at x = 0. A y -intercept of 3 means we have the initial conditional of y(0) = 3 \[3=ce ^{3(0)^{15}}\rightarrow 3=c\]Particular Solution is then: \[y=3e ^{3x ^{15}} \]

OpenStudy (anonymous):

yeah that's what I thought

OpenStudy (anonymous):

Thank you all!